OpenStudy (roberts.spurs19):

Find the cartesian form of r(1 + sin θ ) =a

OpenStudy (rational):

\(x=r\cos\theta\) \(y=r\sin\theta\) \(x^2+y^2=r^2\)

OpenStudy (rational):

\[r(1+\sin\theta)=a\] \[r+r\sin\theta=a\] \[r+y=a\] \[r=a-y\] square both sides and replace \(r^2\) by \(x^2+y^2\)

OpenStudy (roberts.spurs19):

Ah ok thank you! So I think I end up with \[y = \sqrt{a ^{2}-2x ^{2}}\]

OpenStudy (rational):

doesn't look correct how did u get that ?

OpenStudy (roberts.spurs19):

i did \[r^2 = a^2 - x^2\] \[x^2 +y^2 = a^2 -x^2\] \[y^2 = a^2 -2x^2\] which then gave me y = \[\sqrt{a^2 - x^2}\]

OpenStudy (rational):

Lets take it step by step you have : \[r=a-y\] squaring both sides gives you \[r^2=(a-y)^2\] right ?

OpenStudy (roberts.spurs19):

thank you! yep

OpenStudy (rational):

good, simply replace left hand side by \(x^2+y^2\). you get \[x^2+y^2=(a-y)^2\]

OpenStudy (rational):

expand right hand side and cancel y^2 \[x^2+y^2=a^2-2ay+y^2\] \[x^2=a^2-2ay\]

OpenStudy (rational):

which is same as \[x^2+2ay=a^2\]

OpenStudy (roberts.spurs19):

ok then to make y the subject it would be \[y= \frac{ a^2 - x^2} { 2a }\]

OpenStudy (rational):

looks good!

OpenStudy (roberts.spurs19):

Thank you so much for your help!!

OpenStudy (rational):

yw